2. Find the eigenvalues of A= [ 10 -2 1 -5 -2 2 3 -5 1 3 5 and verify that the eigenvectors that mutually orthogonal.
Answer #1

Given matix is A= -5 1 1 10 -2 22 -5. 3 5 we have to find eigenvalue of A for this consider det (AndI) so a I lood | -2 5 -2d3 + 17d² _42dzo & d3 - 178+42dzo A d ( d²-174+42) = 0 5 dl Cd-3) Cd-14)) == 0 d (-3)Cd-14)=0 B Z tod= 0,3, 14 inchezig arreR → Gori | -V न | 0 | -2 2 3 | 0 | |-5 350 २. > R+224 →ps+5 । । - न । 075 20 ___० 2 5 /- | 0 K2- (F) | ना 08 | 0 0 ना । 514 |from above matrix we have vo N, - + z = 0 a Vz+ 5 5 V3=0 a t Here so is free vorable -- put Vart S = + = h tt & Vo = { vz = -10-3 -2 as -2 -5 2 3 3 3 3 5- TV 1 vel lao 0 000 17_-2 -5 TviTTO -2 - 1 3 Vol و اک۔ الافعال we solve this system to find eigeRe 11 - (-2) 22 -247 -5/+10] R₂ (7) 3 L -27 -57% R₂ -To R3-R2 -217 -5/#lo 1 * R4 – (-43) B Lowo OL - TO Looool From above vi: Ve= vist veit v T t T Th | Val P ati I vå J Lt J Lut put t=1 Vaa l is eigenvector corrosponding to eigenvalue d2=3 T InowI 10-14 -2 -5 -2 -5 Tvill 2-143 F 31 5-14 / Lv 1 -2 -12 3 || Val we solve this system of linear egn to find eigenector v = |vIL 1.6 V į 514 the 000 R → R3 (4) T / 514 lol lo 712 o Loooo ke R - R2 10 312 10 o L te lo from above V + V = 0 & vind is fput tot Lis eigenvector nosociated F with eigenvalye. Aos = 16 ale 7 V= 1-5/4 LI are A eigenvector Of Herre Vivoa 15 Szetta-4L V. v3=0 vo Vezo Voo V3=0 . Observe that Leve U₂=0 Li v e and va to each others are mautually cothogonal